myininaya (myininaya):

Say you have the recurrence relation \[a_n-2 a_{n-1}=e^{n} \text{ with } a_0=\frac{2e-2}{e-2} \]... And we want to find a explicit form for the relation. The only way I know of is...

myininaya (myininaya):

\[\text{ charecterisitc equation is } r-2=0 \\ r=2 \\ \text{ so } a_{n_h}=c(2)^n \\ \text{ now particular solution } a_{n_{p}} \\ \text{ so } a_{n_p}=k e^{n} \\ a_{{n-1}_p}=ke^{n-1} \\ ke^{n}-2ke^{n-1}=e^{n} \\ k-2ke^{-1}=1 \\ k=\frac{1}{1-2e^{-1}}=\frac{e}{e-2} \\ \text{ so the solution is } \\ a_n=c (2)^n+\frac{e}{e-2} e^{n} \\ \text{ where we need to apply the } a_0=\frac{2e-2}{e-2} \text{ thing } \\ a_0=c(2)^0+\frac{e}{e-2}e^0 \\ \frac{2e-2}{e-2}=c+\frac{e}{e-2} \\ \frac{2e-2-e}{e-2}=c \\ c=\frac{e-2}{e-2}=1 \\ \text{ so the solution is } a_n=(2)^n+\frac{e}{e-2}e^{n}\]

myininaya (myininaya):

my question is what are other ways of finding this...

myininaya (myininaya):

or names of the ways

myininaya (myininaya):

by the way that one thing above meant homegenous solution the whole \[a_{n_h}\]

OpenStudy (amistre64):

i believe ganeshia does it by use of a generating function ... but its all hocus pocus to me

myininaya (myininaya):

hmmm... that is going to take a little reading if I figure it out, I will post the way by generating function thingy maybe I'm not a good googler, but it is hard for me to search what are the ways to find explicit solutions for recurrence relations

OpenStudy (amistre64):

pattern recognition if possible? a{n} -2a{n-1} = e^n a{n} = 2a{n-1} + e^n ------------------- a0 = k a1 = 2k + e^1 a2 = 2(2k+e^1) + e^2 = 4k + 2e^1 + e^2 = 4k + 2e^1 + e^2 a3 = 8k + 4e^1 + 2e^2 + e^3 a4 = 2^4 k + 2^3 e^1 + 2^2 e^2 + 2^1 e^3 + 2^0 e^4 an = 2^n k + (A)

OpenStudy (amistre64):

\[A=\sum_{k=1}^n 2^{n-k}e^{k}\] not sure how thats useful yet

OpenStudy (amistre64):

2^3 e * e/2 = 2^2 e^2 2^2 e^2 * e/2 = 2^1 e^3 its a geometric sum isnt it

myininaya (myininaya):

\[\sum_{n=0}^{\infty} a_n x^{n}=a(x) \\ \text{ so we have } \\ a_n-2 a_{n-1}=e^{n} \\ \text{ multiply } x^{n} \text{ on both sides } \\ a_n x^{n}-2a_{n-1}x^n =e^{n} x^{n} \\ \text{ now...} \\ \sum_{n=1}^{\infty} (a_n x^{n}-2a_{n-1}x^{n})=\sum_{n=1}^{\infty} e^n x^{n} \\ \sum_{n=1}^{\infty} a_n x^{n} - 2 \sum_{n=1}^{\infty} a_{n-1} x^{n} =\sum_{n=1}^{\infty} e^{n} x^{n} \\ \\ a(x)=\sum_{n=0}^{\infty}a_n x^{n} \\ \sum_{n=1}^{\infty} a_n x^{n}=a(x)-a_0=a(x)-\frac{2e-2}{e-2} \\ 2 \sum_{n=1}^{\infty} a_{n-1}x^{n}=2x \sum_{n=1}^{\infty} a_{n-1}x^{n-1}=2x(a(x)-a_0)=2x(a(x)-\frac{2e-2}{e-2}) \\ \] so we have the equation: \[a(x)-\frac{2e-2}{e-2}-2x(a(x)-\frac{2e-2}{e-2})= \sum_{n=1}^{\infty}e^{n} x^{n} \\ a(x)(1-2x)-\frac{2e-2}{e-2}+2x \frac{2e-2}{e-2}= \sum_{n=1}^{\infty} e^{n} x^{n}\] I could go ahead and solve for a(x) but I'm not sure what to do what the sum thingy on the right hand side

OpenStudy (amistre64):

would you want to solve the homogenous first? or is that redundant?

myininaya (myininaya):

for this method I'm not sure yet

myininaya (myininaya):

so anyways you were trying to do a pattern thingy...

myininaya (myininaya):

\[a_n-2a_{n-1}=e^{n} \\ a_0=\frac{2e-2}{e-2} \\ a_1-2a_0=e^{1} \\ a_1-\frac{4e-4}{e-2}=e \\ a_1=e+\frac{4e-4}{e-2}=\frac{e(e-2)+4e-4}{e-2}=\frac{e^2+2e-4}{e-2} \\ a_2-2a_1=e^{2} \\ a_2=e^2+\frac{2e^2+4e-8}{e-2}=\frac{e^3-2e^2+2e^2+4e-8}{e-2} \\ a_2=\frac{e^3+4e-8}{e-2} \\ a_3=e^{3}+\frac{2e^3+8e-16}{e-2}=\frac{e^4-2e^3+2e^3+8e-16}{e-2}=\frac{e^4+8e-16}{e-2} \\ \text{hmmm.... } \\ a_n=\frac{e^{n+1}+2^{n}-2^{n+1}}{e-2}\]

myininaya (myininaya):

one sec checking...

myininaya (myininaya):

oops missed that e thingy

myininaya (myininaya):

\[a_n=\frac{e^{n+1}+2^n e-2^{n+1}}{e-2}\]

myininaya (myininaya):

yep that is the exact same result the pattern thingy worked well here

myininaya (myininaya):

usually it is harder to see the pattern

OpenStudy (amistre64):

the generating function might be the most applicable :)

myininaya (myininaya):

I couldn't get that way to work

myininaya (myininaya):

because I don't know what to do with the sum(x^n*e^n)

myininaya (myininaya):

actually as n goes to infinity that series diverges

myininaya (myininaya):

unless x is in (-1/e,1/e)

OpenStudy (rational):

The recurrence relation \(a_n-la_{n-1} = f(n)\) has the general solution : \(a_n = l^{n}a_0 + \sum\limits_{i=1}^nl^{n-i}f(n)\)

myininaya (myininaya):

\[a(x)-\frac{2e-2}{e-2}-2x(a(x)-\frac{2e-2}{e-2})= \sum_{n=1}^{\infty}e^{n} x^{n} \\ a(x)(1-2x)-\frac{2e-2}{e-2}+2x \frac{2e-2}{e-2}= \sum_{n=1}^{\infty} e^{n} x^{n}\] \[a(x)(1-2x)-\frac{2e-2}{e-2}(1-2x)=-\frac{ex}{ex-1} \\ \text{ where } x \in (\frac{-1}{e},\frac{1}{e}) \\ a(x)=\frac{-ex}{(ex-1)(1-2x)}+\frac{2e-2}{e-2}\]

myininaya (myininaya):

\[\frac{-ex}{(ex-1)(1-2x)}=\frac{A}{ex-1}+\frac{B}{1-2x} \\ -ex=A-2Ax+Bex-B \\ -ex=A-B+x(-2A+Be) \\ A-B=0 \\ A=B \\ -2B+Be=-e \\ B(-2+e)=-e \\ B=\frac{-e}{-2+e}\]

myininaya (myininaya):

\[a(x)=\frac{e}{e-2} \cdot \frac{1}{ex-1}+\frac{e}{e-2} \cdot \frac{1}{1-2x}+\frac{2e-2}{e-2}\]

myininaya (myininaya):

I have to play a little more with this later I have to go to a thingy with some thingy

OpenStudy (anonymous):

I'm a big fan of using GFs whenever possible. Here's a question I answered recently with the method, maybe it'll give you a good idea of how it works? http://openstudy.com/study#/updates/5600180fe4b0ed58e276f560

myininaya (myininaya):

I just noticed I made a mistake above: \[a(x)=\frac{e}{2-e} \cdot \frac{1}{ex-1}+\frac{e}{2-e} \cdot \frac{1}{1-2x}+\frac{2e-2}{e-2} \\ a(x)=\frac{e}{2-e} \cdot \sum_{n=0}^{\infty}-e^nx^n +\frac{e}{2-e} \cdot \sum_{n=0}^{\infty} 2^{n} x^{n} +\frac{2e-2}{e-2} \\ \sum_{n=0}^{\infty} a_n x^{n}=\frac{e}{2-e} \cdot \sum_{n=0}^{\infty} -e^{n} x^n+\frac{e}{2-e} \sum_{n=0}^{\infty} 2^n x^n +\frac{2e-2}{e-2} \\ \text{ somehow this means } \\ a_n=\frac{e}{2-e}(-e^n)+\frac{e}{2-e}2^{n}+\frac{2e-2}{e-2} \\\] I think I did something wrong I will go back over it after reading those links you have posted

myininaya (myininaya):

\[a_n-2 a_{n-1}=e^n \text{ with } a_0=\frac{2e-2}{e-2} \\ \sum_{n=1}^{\infty}(a_n-2 a_{n-1})x_n=\sum_{n=1}^\infty e^{n} x^n \\ \\ \sum_{n=1}^{\infty} (ex)^n \text{ assume } |e x|<1 \implies |x|<\frac{1}{e} \\ \text{ then } \sum_{n=1}^{\infty} (ex)^n=\frac{ex}{1-ex} \\ \text{ Let } a(x)=\sum_{n=0}^{\infty} a_n x^n \\ \sum_{n=1}^{\infty} a_n x^{n}=a(x)-a_0=a(x)-\frac{2e-2}{e-2} \\ -2 \sum_{n=1}^{\infty} a_{n-1}x^{n}= -2 x \sum_{n=1}^{\infty} a_{n-1}x^{n-1}=-2x[ \sum_{k=0}^\infty a_k x^k]=-2x[a(x)]=-2x a(x) \] so I think I see my mistake... \[a(x)-\frac{2e-2}{e-2}-2 x a(x)= \frac{ex}{1-ex} \\ a(x)(1-2x)=\frac{ex}{1-ex}+\frac{2e-2}{e-2} \\ \\ a(x)= \frac{ex}{(1-2x)(1-ex)}+\frac{2e-2}{e-2} \cdot \frac{1}{1-2x} \\ a(x)=\frac{e}{2-e} \cdot \frac{1}{1-2x}-\frac{e}{2-e } \cdot \frac{1}{1-ex}+ \cdot \frac{2e-2}{e-2} \cdot \frac{1}{1-2x}\] \[a(x)=\frac{e}{2-e} \sum_{n=0}^{\infty} 2^n x^n -\frac{e}{2-e} \sum_{n=0}^{\infty} e^n x^n +\frac{2e-2}{e-2} \cdot \sum_{n=0}^{\infty} 2^n x^n \\ a_n=\frac{e}{2-e} 2^{n} -\frac{e}{2-e}( e^n) +\frac{2e-2}{e-2} 2^n \\ a_n =2^n \cdot (\frac{2e-2-e}{e-2})+\frac{e e^n}{e-2 } \\ a_n=2^n + \frac{e^{n+1}}{e-2}\] got it! definitely made a mistake before.. ---partial fraction break down below for that one term: \[\frac{ex}{(1-2x)(1-ex)}=\frac{A}{1-2x}+\frac{B}{1-ex} \\ ex=(A+B)+x(-Ae-2B) \\ A=-B \\ -Ae-2B=e \\ -Ae+2A=e \\ A=\frac{e}{2-e} \\ B=\frac{-e}{2-e}\]

myininaya (myininaya):

I would like to see other strategies if there are any others but I still kinda prefer that homogeneous/particular solution way.

OpenStudy (kainui):

This looks like a scary mess :O

OpenStudy (freckles):

but solving fun scary things make you seem more intimidating and isn't that the whole point